OpenStudy (anonymous):

Find the area of the region that is bounded by the given curve and lies in the specified sector

OpenStudy (anonymous):

\[r=\sqrt \theta , 0\le \theta \le \frac{ \pi }{ 4 }\]

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{ \pi }{ 4 }}\frac{ 1 }{ 2 } \theta d \theta\]

OpenStudy (ankitshaw):

its simple integration .. give it a try..

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\frac{( \frac{ \pi }{ 4 } )^3}{ 3 }\]

OpenStudy (anonymous):

\[\frac{ \frac{ \pi }{ 64 } }{ 6 }\]

OpenStudy (anonymous):

the answer is \[\frac{ \pi^2 }{ 64 }\]

OpenStudy (anonymous):

im just not getting that

OpenStudy (anonymous):

can u help me i'm begging u i dont want to fail please

OpenStudy (anonymous):

please

OpenStudy (anonymous):

with what?

OpenStudy (anonymous):

this

OpenStudy (anonymous):

I am the one that posted this question hun. I need help too

OpenStudy (anonymous):

OK me to and i have no idea on how to get the help?

OpenStudy (anonymous):

just hang tight, maybe someone will come through for us

OpenStudy (anonymous):

i hope so

OpenStudy (anonymous):

on the urge for crying

OpenStudy (ankitshaw):

you have made some mistake in performing integration

OpenStudy (anonymous):

ok...?

OpenStudy (ankitshaw):

\[\int\limits \theta d \theta = \theta^2 /2\]

OpenStudy (ankitshaw):

i hope this help now try to do it correctly. with your question.

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{ \pi }{ 4 }}\frac{ 1 }{ 2 }(\theta)^2d \theta= \frac{ 1 }{ 2 }\int\limits_{0}^{\frac{ \pi }{ 4 }}\theta d \theta = \frac{ 1 }{ 2 }(\frac{ \theta^2 }{ 2 })\]\[= \frac{ 1 }{ 2 }(\frac{ (\frac{ \pi }{ 4 })^2 }{ 2 })= \frac{ 1 }{ 2 }(\frac{ \pi^2 }{ 32 })= \frac{ \pi^2 }{ 64 }\]

OpenStudy (anonymous):

is this correct?

OpenStudy (ankitshaw):

yes

OpenStudy (anonymous):

thank you so much!

OpenStudy (ankitshaw):

welcome..

OpenStudy (anonymous):

@home_work check out my steps here hun